OpenStudy (anonymous):

Find dy/dx by implicit differentiation. (sin πx + cos πy)9 = 34

OpenStudy (bahrom7893):

hey lostdove..

OpenStudy (anonymous):

hey how r u

OpenStudy (bahrom7893):

So more differentiation and calculus http://www.twiddla.com/579399

OpenStudy (anonymous):

can u please help me with this problem ive been trying to do it forever lol

OpenStudy (anonymous):

is that to the power of 9?

OpenStudy (anonymous):

\[(\sin(\pi x)+\cos(\pi y))^9=34\]?

OpenStudy (anonymous):

dang did it again.

OpenStudy (anonymous):

\[9(\sin(\pi x)+ \cos(\pi y)^8\times (\pi \cos(\pi x)-\pi \sin(\pi y)y')=0\]

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

power of 9